# On Efficient Variants of Segment Anything Model: A Survey

### v3

@@ 4.1 Datasets and Metrics
We utilize two fundamental datasets, COCO 2017 [219] and LVIS v1 [220], and two more challenging benchmarks, SGinW [^1] https://eval.ai/web/challenges/challenge-page/1931/overview and UVO v1.0 [221], for comprehensive evaluation. COCO is a large-scale dataset for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS.
Segmentation in the Wild (SGinW) benchmark consists of 25 different segmentation datasets, spanning a wide range of real-world domains, from common objects like fruits and animals to rare concepts like rail and brain tumor[222]. Unidentified Video Objects (UVO) is a challenging benchmark for dense, open-world segmentation in videos, with numerous ”unseen” objects not belonging to any of the COCO categories or the LVIS categories. It is worth noting that, for UVO benchmark, we only leverage the FrameSet for image-level evaluation, which includes 8124 frames sampled from 2708 videos with 65588 annotated instances.
To evaluate efficiency, we first test several soft indicators like #$\#Params$, $FLOPs$, and $Memory\ Usage$. And we further calculate the Efficient Error Rate ($EER$) for a more comprehensive evaluation, as outlined in [55], to measure how much the variants are more efficient in comparison with the vanilla SAM-H. Specifically, $EER$ is calculated by averaging the ratios of a set of efficient metrics between evaluated models and the reference model. The definition of $EER$ is illustrated as follows,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. Consequently, tested models with higher efficiency will achieve lower $EER$. In our evaluation, we select {#$\#Params$, $FLOPs$} as metrics and set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of these models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
[…]
@@
**footnotetext: The results of FastSAM [41] are collected from [46] and others are evaluated under our framework.
@@ 4.2 Efficiency Comparison
We first report the efficiency results of SAM and its variants. With the picture groceries.jpg [^2] https://github.com/facebookresearch/segment-anything/blob/main/notebooks/images/groceries.jpg used in SAM’s official example as input, we utilize a bounding box as prompt and evaluate models’ #$\#Params$ and $FLOPs$ by leveraging the tool calflops [^3] https://github.com/MrYxJ/calculate-flops.pytorch. We then calculate the $EER$ for further comparison. The results are illustrated in Tab. 3. Among the efficient variants, we observe that EdgeSAM has the lowest number of parameters, $FLOPs$ and consequently the $EER$, while EfficientViT-SAM-XL1 takes the highest numbers whose $EER$ is even 6.51% more than SAM-B. When compared to the heaviest SAM-H, all variants present obvious reduction in both model size and calculation.
[…]
@@
Table 5: Quantitative results of inference latency (ms) of SegEvery task with grid sampling in different scales or other efficient strategies. More details are reviewed in Section 3.2.
@@ 4.3 Accuracy Comparison
In this subsection, we report the accuracy results of SAM and its variants on SegAny task (with point/box prompts) and instance segmentation task, respectively. We follow the evaluation framework in [122] and adapt it by introducing other evaluation modules to conduct a unified evaluation on variants.
Table 6: Quantitative results of the accuracy of SegAny task (mIoU) on COCO and LVIS with points and boxes as prompts. For evaluation with points prompts, we select the center point of the ground truth bounding box ($pt_{1}$), and one or three randomly sampled points from ground truth masks (1*$pt_{2}$, 3*$pt_{2}$). For evaluation with boxes prompts, we adopt the ground truth bounding box ($box_{1}$), and the tightest bounding box calculated by the ground truth mask ($box_{2}$).
@@
Table 7: Quantitative results of instance segmentation on COCO and LVIS with ViTDet [223] as object detector.
@@
To evaluate on the SegAny task, we adopt two types of points as prompts: 1) the center point of the ground truth bounding box, and 2) random points uniformly sampled from ground truth masks, following the setting in [109]. We evaluate the variants on COCO and LVIS, and the mean intersection over union (mIoU) is reported in Tab. 6. When prompted with the center point, SAM2-B+ and EfficientViT-SAM-XL1 achieve the highest mIoU of 54.3% on COCO dataset, surpassing SAM-H with 53.6% mIoU, while on the LVIS dataset, SAMfast-H, with mIoU of 53.6%, exhibits the best performance among variants. Under the setting of random point prompts, it is demonstrated that EfficientViT-SAM-XL1 performs better than SAM-H on both datasets when prompted with 3 points, with the increase of 2.7% and 0.7%, respectively. From the perspective of datasets, we observe that the results on LVIS are generally lower than those on COCO.
Moreover, we also evaluate the accuracy on SegAny task with two types of box prompts: 1) The ground truth bounding box and 2) The tightest bounding box corresponding to the ground truth mask, instructed by the experiments in [109, 122]. We reported the results of mIoU on COCO and LVIS which are illustrated in Tab. 6. We observe that EfficientViT-SAM-XL1 demonstrate the highest accuracy in every box-prompted setting, which exceeds SAM-H by 2.5%, 2.0%, 1.9% and 0.6%, respectively. SAMfast-H and EfficientViT-SAM-L0 also present performance close to SAM-H in the segmentation tasks with box prompts.
Table 8: Quantitative results of instance segmentation on COCO with YOLOv8 [169] or GrounddingDINO [224] as object detector.
@@
Table 9: Quantitative results of instance segmentation on COCO with Detic [225] or H-Deformable-DETR [226] as the object detector.
@@
For instance segmentation tasks, we adopt the ViTDet [223], YOLOv8 [169], GrounddingDINO [224], Detic [225] and H-Deformable-DETR [226] with Swin-L [227] as object detectors to generate bounding boxes of latent objects, with reference to the experiments conducted in [113, 46, 122]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 7, 8 and 9. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors. Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
To further evaluate variants’ segmentation ability in more challenging scenarios, we conduct instance segmentation on SGinW and UVO benchmark, equipped with GrounddingDINO for box prompts generation by following [104]. Results are illustrated in Tab. 10 and Tab. 11, respectively. On SGinW benchmark, we observe that the rewritten SAMfast-H achieves the highest mean AP over 25 datasets, with only 0.1% decrease compared to SAM-H, followed by EfficientViT-SAM-XL1 and EfficientSAM-S with 0.4% and 0.8% more drop respectively. From the perspective of dataset, the segmentation results on House-Parts, Brain-Tumor and Rail are significantly lower than others, indicating them more challenging datasets. On UVO benchmark, we notice that EfficientViT-SAM-XL1 demonstrates the best performance with AP of 31%, exceeding SAM-H by 1.1%, while SAM2.1-B+ also surpasses original SAM by 1.0%.

### v4

@@ 4.1 Datasets and Metrics
We utilize two fundamental datasets, COCO 2017 [lin2014microsoft] and LVIS v1 [gupta2019lvis], and two more challenging benchmarks, SGinW [^1] https://eval.ai/web/challenges/challenge-page/1931/overview and UVO v1.0 [wang2021unidentified], for comprehensive evaluation. COCO is a large-scale dataset for object detection, segmentation, and captioning, containing 330K images and 1.5 million object instances across 80 object categories. LVIS is tailored for large vocabulary instance segmentation, featuring over 2 million high-quality segmentation masks across more than 1200 categories in 164K images. For our evaluation, we use the validation sets of both datasets, which include 36,781 instances in 5,000 images for COCO and 244,707 instances in 19,809 images for LVIS.
Segmentation in the Wild (SGinW) benchmark consists of 25 different segmentation datasets, spanning a wide range of real-world domains, from common objects like fruits and animals to rare concepts like rail and brain tumor[zou2023generalized]. Unidentified Video Objects (UVO) is a challenging benchmark for dense, open-world segmentation in videos, with numerous ”unseen” objects not belonging to any of the COCO categories or the LVIS categories. It is worth noting that, for UVO benchmark, we only leverage the FrameSet for image-level evaluation, which includes 8124 frames sampled from 2708 videos with 65588 annotated instances.
To evaluate efficiency, we first test several soft indicators like #Params\#Params, FLOPsFLOPs, and MemoryUsageMemory\ Usage. And we further calculate the Efficient Error Rate ($EER$) for a more comprehensive evaluation, as outlined in [papa2024survey], to measure how much the variants are more efficient in comparison with the vanilla SAM-H. Specifically, $EER$ is calculated by averaging the ratios of a set of efficient metrics between evaluated models and the reference model. The definition of $EER$ is illustrated as follows,
@@
where $N$ is the number of metrics and $M_{i}$, $R_{i}$ refer to the i-th metric of the tested model and the reference one, respectively. Consequently, tested models with higher efficiency will achieve lower $EER$. In our evaluation, we select {#Params\#Params, FLOPsFLOPs} as metrics and set the reference model to be SAM-H. In addition to these metrics, we also report the runtime and throughput of these models. For accuracy evaluation, we use mean intersection over union ($mIoU$) for the SegAny task and mean average precision ($AP$) for instance segmentation.
Table 3: Quantitative results of $\textbf{Params (M)}$, FLOPs (G) and EER. EEREER denotes the Efficient Error Rate proposed in [papa2024survey] and is calculated by Eq. (11), which averaging the ratios of {#Params, FLOPs} between variants and the vanilla SAM-H. The lower value of EEREER represent higher efficiency of the tested variant.**footnotetext: The results of FastSAM [zhao2023fast] are collected from [zhou2023edgesam] and others are evaluated under our framework.
@@ 4.2 Efficiency Comparison
We first report the efficiency results of SAM and its variants. With the picture groceries.jpg [^2] https://github.com/facebookresearch/segment-anything/blob/main/notebooks/images/groceries.jpg used in SAM’s official example as input, we utilize a bounding box as prompt and evaluate models’ #Params\#Params and $FLOPs$ by leveraging the tool calflops [^3] https://github.com/MrYxJ/calculate-flops.pytorch. We then calculate the $EER$ for further comparison. The results are illustrated in Tab. 3. Among the efficient variants, we observe that EdgeSAM has the lowest number of parameters, $FLOPs$ and consequently the $EER$, while EfficientViT-SAM-XL1 takes the highest numbers whose $EER$ is even 6.51% more than SAM-B. When compared to the heaviest SAM-H, all variants present obvious reduction in both model size and calculation.
We have also measured the models’ inference time in both SegAny and SegEvery modes, using images from the COCO validation set as evaluation data. For the SegAny task, each image is prompted with 50 fixed bounding boxes. We record the cumulative time for every 10 boxes and illustrate the results through a curve chart (shown in Fig. 17). Based on this, we calculate the average time required to process one image with one box prompt and report it as the inference latency for the SegAny task. We conduct this evaluation on CPU-only environment, a series of single-GPU environments, and the edge device JetsonNano 4GBJetson\ Nano\ 4GB. We simultaneously report the approximation of GPU memory usage when running these models. Additionally, we test the throughput of each variant of conducting SegAny task on the COCO validation set, using ground truth bounding boxes as prompts. The results are summarized in Tab. 4. Our findings reveal that EfficientViT-SAM-L0 achieves the shortest inference time for the SegAny task, offering nearly 30x acceleration on GPU and almost 50x acceleration on CPU compared to the heaviest model, SAM-H. EdgeSAM also shows impressive performance with low latencies of 259 ms on CPU and 713 ms on the edge device, while NanoSAM achieves the lowest runtime on $JetsonNano$ and comes close to the best results on all GPU environments. In the throughput test, NanoSAM leads with 27.9 images processed per second. Two other variants, EfficientSAM-Ti and EfficientViT-SAM-L0, also demonstrate strong throughput, each exceeding 20 images per second.
@@
Table 5: Quantitative results of inference latency (ms) of SegEvery task with grid sampling in different scales or other efficient strategies. More details are reviewed in Section 3.2.**footnotetext: OOM denotes out of memory.
@@ 4.3 Accuracy Comparison
In this subsection, we report the accuracy results of SAM and its variants on SegAny task (with point/box prompts) and instance segmentation task, respectively. We follow the evaluation framework in [zhang2024efficientvit] and adapt it by introducing other evaluation modules to conduct a unified evaluation on variants.
[…]
@@
Table 7: Quantitative results of instance segmentation on COCO and LVIS with ViTDet [li2022exploring] as object detector.
@@
To evaluate on the SegAny task, we adopt two types of points as prompts: 1) the center point of the ground truth bounding box, and 2) random points uniformly sampled from ground truth masks, following the setting in [xiong2024efficientsam]. We evaluate the variants on COCO and LVIS, and the mean intersection over union (mIoU) is reported in Tab. 6. When prompted with the center point, SAM2-B+ and EfficientViT-SAM-XL1 achieve the highest mIoU of 54.3% on COCO dataset, surpassing SAM-H with 53.6% mIoU, while on the LVIS dataset, SAMfast-H, with mIoU of 53.6%, exhibits the best performance among variants. Under the setting of random point prompts, it is demonstrated that EfficientViT-SAM-XL1 performs better than SAM-H on both datasets when prompted with 3 points, with the increase of 2.7% and 0.7%, respectively. From the perspective of datasets, we observe that the results on LVIS are generally lower than those on COCO.
Moreover, we also evaluate the accuracy on SegAny task with two types of box prompts: 1) The ground truth bounding box and 2) The tightest bounding box corresponding to the ground truth mask, instructed by the experiments in [xiong2024efficientsam, zhang2024efficientvit]. We reported the results of mIoU on COCO and LVIS which are illustrated in Tab. 6. We observe that EfficientViT-SAM-XL1 demonstrate the highest accuracy in every box-prompted setting, which exceeds SAM-H by 2.5%, 2.0%, 1.9% and 0.6%, respectively. SAMfast-H and EfficientViT-SAM-L0 also present performance close to SAM-H in the segmentation tasks with box prompts.
Table 8: Quantitative results of instance segmentation on COCO with YOLOv8 [yolov8] or GrounddingDINO [liu2023grounding] as object detector.
@@
Table 9: Quantitative results of instance segmentation on COCO with Detic [zhou2022detecting] or H-Deformable-DETR [jia2023detrs] as the object detector.
@@
For instance segmentation tasks, we adopt the ViTDet [li2022exploring], YOLOv8 [yolov8], GrounddingDINO [liu2023grounding], Detic [zhou2022detecting] and H-Deformable-DETR [jia2023detrs] with Swin-L [liu2021swin] as object detectors to generate bounding boxes of latent objects, with reference to the experiments conducted in [wang2023repvit, zhou2023edgesam, zhang2024efficientvit]. We evaluate the average precision (AP) of all objects, as well as AP of small, medium, and large objects. The results are reported in Tab. 7, 8 and 9. Similar to the previous results, we find that on COCO dataset EfficientViT-SAM-XL1 always presents the highest AP with any of the detectors. Under the setting of equipping ViTDet as detector and testing on LVIS dataset, SAMfast-H surpass all other variants with AP of 44.5%.
To further evaluate variants’ segmentation ability in more challenging scenarios, we conduct instance segmentation on SGinW and UVO benchmark, equipped with GrounddingDINO for box prompts generation by following [ke2024segment]. Results are illustrated in Tab. 10 and Tab. 11, respectively. On SGinW benchmark, we observe that the rewritten SAMfast-H achieves the highest mean AP over 25 datasets, with only 0.1% decrease compared to SAM-H, followed by EfficientViT-SAM-XL1 and EfficientSAM-S with 0.4% and 0.8% more drop respectively. From the perspective of dataset, the segmentation results on House-Parts, Brain-Tumor and Rail are significantly lower than others, indicating them more challenging datasets. On UVO benchmark, we notice that EfficientViT-SAM-XL1 demonstrates the best performance with AP of 31%, exceeding SAM-H by 1.1%, while SAM2.1-B+ also surpasses original SAM by 1.0%.